What is the average KE of an air molecule in a room? (T = 21 degrees)
Answer #1

To determine the average kinetic energy of gas molecules we need to know the temperature of the gas, the universal gas constant (R), and Avogadro's number (Na).

K = (3/2)*R*T/Na

where K is average kinetic energy of air molecule.

R is gas constant which is equal to (8.314 J/mol.k)

T is temperature in Kelvin

Na is the Avogadro's Number (6.022 * 10^23 atoms/mol)

using the values,

K = (1.5) * 8.314 * 294/(6.022 * 10^23)

K = 6.08 * 10^-21 J/atom

Thus the average kinetic energy of an air molecule in 21° C temperature is 6.08 * 10^-21 J/atom.